a force applied on a body of1.2kg for 0.2 s and the body travelled a distance of 2m in the next round. Find the force applied on th body.
1
Expert's answer
2019-07-24T15:49:57-0400

The Newton's second law states

"F=ma"


The distance travelled by body

"d=\frac{at^2}{2}"

So

"F=m\times \frac{2d}{t^2}"

"=1.2\:\rm{kg}\times \frac{2\times 2\:\rm{m}}{(0.2\:\rm{s})^2}=120\:\rm{N}"